You conduct a mark and recapture experiment to measure the size of a population of birds in your area. In your initial capture, you tag 173 birds. In your second capture, you observe 212 birds and find that 15 of them are tagged. What is the estimated size of the population? (Note: round to the nearest individual, but only at the end of the calculation.)
Answer #1

Let the size of the population is N. Then,

The proportion of tagged and total in both population and sample should be same. Hence,

N/173 = 212/15

Therefore,

N = (212/15)*173

N = 2445